Jake is younger than his brother by 10 years. Five years ago the sum of their ages was 60. Find their present age.
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Let J be jake, b be his brother:
b=j+10
(j+10)-5+j-5=60
2j=60
j=30
j+10=40
Jake is 30, his brother is 40!!!!!!!!!!!!!!!

Answer by ikleyn(52943)

.
Jake is younger than his brother by 10 years. Five years ago the sum of their ages was 60. Find their present age.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

For detailed solutions of age problems see the lessons
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.